SecuriCorp operates a fleet of armored cars that make scheduled pickups and deliveries in the Los Angeles area. The company is implementing an activity-based costing system that has four activity cost pools: Travel, Pickup and Delivery, Customer Service, and Other. The activity measures are miles for the Travel cost pool, number of pickups and deliveries for the Pickup and Delivery cost pool, and number of customers for the Customer Service cost pool. The Other cost pool has no activity measure because it is an organization-sustaining activity. The following costs will be assigned using the activity-based costing system:

Driver and guard wages $ 1,040,000
Vehicle operating expense 470,000
Vehicle depreciation 350,000
Customer representative salaries and expenses 380,000
Office expenses 240,000
Administrative expenses 540,000
Total cost $ 3,020,000

The distribution of resource consumption across the activity cost pools is as follows:

Travel Pickup
and
Delivery
Customer
Service
Other Totals
Driver and guard wages 50 % 35 % 10 % 5 % 100 %
Vehicle operating expense 70 % 5 % 0 % 25 % 100 %
Vehicle depreciation 60 % 15 % 0 % 25 % 100 %
Customer representative salaries and expenses 0 % 0 % 90 % 10 % 100 %
Office expenses 0 % 20 % 30 % 50 % 100 %
Administrative expenses 0 % 5 % 60 % 35 % 100 %

Required:

Complete the first stage allocations of costs to activity cost pools.
Answer #1
Travel Pickup and Delivery Customer Service Other Total
Driver and guard wages 520000 364000 104000 52000 1040000
Vehicle operating expense 329000 23500 0 117500 470000
Vehicle depreciation 210000 52500 0 87500 350000
Customer representative salaries and expenses 0 0 342000 38000 380000
Office expenses 0 48000 72000 120000 240000
Administrative expenses 0 27000 324000 189000 540000
Total cost 1059000 515000 842000 604000 3020000
Workings:
Travel Pickup and Delivery Customer Service Other
Driver and guard wages =1040000*50% =1040000*35% =1040000*10% =1040000*5%
Vehicle operating expense =470000*70% =470000*5% 0 =470000*25%
Vehicle depreciation =350000*60% =350000*15% 0 =350000*25%
Customer representative salaries and expenses 0 0 =380000*90% =380000*10%
Office expenses 0 =240000*20% =240000*30% =240000*50%
Administrative expenses 0 =540000*5% =540000*60% =540000*35%